# On the robustness of modeling grounded word learning through a child’s egocentric input

### v1

@@ Abstract
What insights can machine learning bring to understanding human language acquisition? Large language and multimodal models have achieved remarkable capabilities, but their reliance on massive training datasets creates a fundamental mismatch with children, who succeed in acquiring language from comparatively limited input. To help bridge this gap, researchers have increasingly trained neural networks using data similar in quantity and quality to children’s input. Taking this approach to the limit, Vong et al., (2024) showed that a multimodal neural network trained on 61 hours of visual and linguistic input extracted from just one child’s developmental experience could acquire word-referent mappings. However, whether this approach’s success reflects the idiosyncrasies of a single child’s experience, or whether it would show consistent and robust learning patterns across multiple children’s experiences was not explored. In this article, we applied automated speech transcription methods to the entirety of the SAYCam dataset, consisting of over 500 hours of video data spread across all three children. Using these automated transcriptions, we generated multi-modal vision-and-language datasets for both training and evaluation, and explored a range of neural network configurations to examine the robustness of simulated word learning. Our findings demonstrate that networks trained on automatically transcribed data from each child can acquire and generalize word-referent mappings across multiple network architectures. These results validate the robustness of multimodal neural networks for grounded word learning, while highlighting the individual differences that emerge in how models learn when trained on each child’s developmental experiences.
@@ 1 Introduction
Recent progress in large language models and multi-modal AI systems have led to fluent language generation and accurate object recognition, among other human-like abilities. These developments have naturally piqued the interest of many cognitive scientists and developmental psychologists, raising the question of whether these computational advances can help explain how children acquire language in the first few years of their lives. Yet, there is a striking asymmetry to reckon with: can machines trained to learn language from billions, even trillions, of word tokens provide insight into how children learn language from mere millions (Frank,, 2023)? Addressing this disparity requires examining whether these language learning mechanisms can succeed under more realistic developmental constraints, combining these computational methods with richer and more ecologically valid datasets from developmental psychology.
@@
Recent efforts towards this goal have examined what can be learned from linguistic (and other sensory) data that is comparable in quantity or kind to what children receive naturally (Huebner et al.,, 2021; Wang et al.,, 2023; Warstadt et al.,, 2024; Qin et al.,, 2023; Vong et al.,, 2024; Hu et al.,, 2024; Zhuang et al.,, 2024). For example, Huebner et al., (2021) examined language learning in neural networks from aggregated child-directed speech corpora (MacWhinney,, 1992), demonstrating that models can acquire syntactic and semantic knowledge in the limited-data regime that is more representative of the kinds of input children receive. Similarly, the BabyLM challenge (Warstadt et al.,, 2024) examined what kinds of linguistic competencies can be learned from training language models on datasets consisting of 10M to 100M words (Hu et al.,, 2024; Charpentier et al.,, 2025). Separately, Zhuang et al., (2024) studied how visual information can improve language learning, especially when using developmentally plausible amounts of data.
@@
Additional recent work has aimed to increase the realism of the training data by using egocentric, head-mounted camera recordings from developing children, such as SAYCam (Sullivan et al.,, 2022) and BabyView (Long et al.,, 2024). These datasets provide a unique glimpse into what children actually see and hear over the course of language learning, providing a testbed for what can be learned by training models from this kind of naturalistic, high-quality input. Researchers have shown that certain kinds of syntax and semantics, as well as word-referent mappings, are all learnable, even under the extreme constraint of only using data collected from a single child (Wang et al.,, 2023; Qin et al.,, 2023; Vong et al.,, 2024; Long et al.,, 2024).
@@
In recent work most related to the current article, Vong et al., (2024) demonstrated that a multimodal neural network trained on vision-and-language data extracted from a single child could acquire word-referent mappings. Specifically, they used a subset of a single child in SAYCam Sullivan et al., (2022), curating a dataset of paired child-directed utterances with corresponding video frames from 61 hours of video. This dataset was used to train a multimodal neural network utilizing two separate encoders for visual and linguistic inputs, combined with a contrastive objective to align embeddings from both modalities into a shared multimodal representation space. The model’s knowledge of word-referent mappings was evaluated on two tasks involving 22 within-distribution categories and 64 out-of-distribution categories, achieving high performance on the former and modest performance on the latter. This work demonstrated that grounded word-referent mappings could be acquired from a subset of one child’s experience from relatively generic neural networks. However, a major limitation is that its findings are reliant on the data from a single child (as only one child’s data had transcripts at the time), raising questions about how robust or generalizable the findings are.
@@
In this article, we examine the robustness of the findings reported in Vong et al., (2024) by utilizing the full 500 hours of video from the SAYCam dataset (Sullivan et al.,, 2022) and training distinct models on each child’s data. The main contributions of our study are:
@@
  • We also explore robustness across different language model configurations and training objectives, finding modest variations in performance, suggesting that dataset characteristics outweigh specific modeling or architectural choices.
@@ 2 Datasets
Our approach requires datasets that closely mirror the visual and linguistic experience from the perspective of developing infants. One dataset that fits this criterion is SAYCam(Sullivan et al.,, 2022), a video dataset capturing the egocentric, longitudinal perspective of 3 different children aged 6 to 32 months, collected via head-mounted cameras. While previous attempts to train language and multimodal models from SAYCam showed promise in answering questions around learnability (Wang et al.,, 2023; Vong et al.,, 2024; Qin et al.,, 2023), their results were limited by the availability of manually transcribed natural language data, primarily to a subset of one of the three children’s data. In this work, we expand the amount of available data for training by a factor of 7 through the use of automated speech transcription tools (Radford et al.,, 2023; Bain et al.,, 2023). This enabled the creation of new training and evaluation datasets for each of the three children in SAYCam, providing an opportunity to replicate our studies of simulated word learning with multiple children’s data.
@@ 2.1 Training Datasets
Transcribing videos. In order to train separate vision-language models on each child’s data, we needed a way to transcribe all of the speech from the raw videos. To efficiently transcribe the long-form audio from 500 hours of video from SAYCam, we used WhisperX (Bain et al.,, 2023), a tool which leverages the transcription capabilities of OpenAI’s Whisper model (Radford et al.,, 2023) (using the v3-large model with the language set to English), and applies further processing for efficient transcription and speaker diarization and word-level timestamps. This provided us with the transcribed text of each separate utterance, the start and end timestamps, and the predicted speaker label. Additional details surrounding the automatic transcription procedure and evaluation for speech transcription accuracy can be found in Zhang et al., (2025).
@@
Due to the inherent challenges of recording audio from a child’s head-mounted camera, the audio quality, especially for child-directed speech is highly variable across videos. Nevertheless, an analysis comparing a subset of manually vs. automatically transcribed speech content showed that it was relatively accurate for adult speech and also sufficiently similar to the previously manually transcribed subset (Zhang et al.,, 2025). Furthermore, although the relative timings for the detected start and end timestamps also showed some discrepancies, this is not necessarily a major issue: Vong et al., (2024) observed that imperfect timing information (in the manual transcriptions) were still sufficient for models to acquire word-referent mappings.
@@
Whisper was not a panacea for annotating the data: it showed relatively poor performance on diarization accuracy (correctly distinguishing between different speakers in a video). Therefore, we did not attempt to distinguish between different speakers (e.g. child-directed versus child-produced speech), and retained all of the transcribed utterances. Another limitation was from hallucinations in the Whisper transcriptions, which resulted in instances of repetitive utterances, [^1] This included filtering repetitive phrases within a single utterance, as well as filtering adjacent utterances with the same content. In both cases, we only kept a single instance of the repeated utterance., requiring them to be filtered out. Additional pre-processing steps included removing all punctuation and lower-casing the transcribed text. Tokenization was performed at the word level using spaCy (Honnibal and Montani,, 2017) and was subsequently used to create separate tokenizers from the transcribed vocabulary of each dataset split (after excluding any words with a frequency of two or fewer). After filtering and pre-processing, this resulted in 266K automatically transcribed utterances across the entire SAYCam dataset.
@@
Training datasets. To create vision-language datasets for training, we used this set of transcribed utterances to extract video frames in their original resolution that were temporally aligned with each utterance. For each detected utterance, we used the detected starting timestamp to extract up to 16 frames from the video at a rate of 3.75 fps, corresponding to 4.27s of video. Each set of extracted of video frames was paired with their corresponding utterance, forming the basis of our vision-language dataset for training. The total amount of available training data from automating the transcription process is roughly 7 times larger than what was previously available for a single baby via manual transcription (Sullivan et al.,, 2022; Vong et al.,, 2024). We split the utterances into five different dataset splits, making sure that each split uses data from at most one child. The five different splits are:
@@
  • S-Whisper-2022: This split is based on the subset of the Whisper utterances from the same set of videos that were previously manually transcribed for baby S, as described in Sullivan et al., (2022).
@@ 2.2 Evaluation Datasets
Figure 1: Evaluation frames from Labeled-S-V2, Labeled-A and Labeled-Y. Here we present three randomly selected evaluation frames from four different evaluation categories (ball, toy book and cereal), with each row indicating frames derived from each of the three children in the SAYCam dataset.
For model evaluation, we primarily focus on image classification tasks, which are based on the looking-while-listening paradigm from early word learning studies (Golinkoff et al.,, 1992). In this evaluation procedure, models are presented with a target word along with four video frames (one belonging to the target category, and three distractors from other randomly sampled categories). The task is to select the video frame corresponding to the target word, where these labeled frames are sourced from SAYCam directly. We test our models on a number of different evaluation datasets:
  • Labeled-S: This first evaluation set consists of 22 different visual categories derived from baby S’s egocentric video data, covering the most common kinds of objects and places both seen and spoken about (Orhan et al.,, 2020; Vong et al.,, 2024). For each category, there are 100 separate evaluation trials for a total of 2200 evaluation trials. The full set of categories can be found in Appendix C.
  • Labeled-S-V2, Labeled-A and Labeled-Y: In addition to Labeled-S, we create additional evaluation datasets based on the datasets from each child, using a mixture of automatic and manual filtering (see Appendix A for additional details). For each category, there are 100 separate evaluation trials, for a total of 4100 evaluation trials. The frames used for evaluation were (primarily) derived from the test splits from each baby. [^2] For Labeled-S-V2, due to some differences in dataset splitting, there was some partial overlap between the videos used for extracting frames for training and evaluation. These 41 categories cover common visual concepts encountered by each child in SAYCam, including body parts (arm, hair, knee), clothing (shirt, jeans, shoe), living spaces (kitchen, porch, table), objects (ball, bottle, boy) and food (banana, cereal). A limited set of evaluation frames are shown in Figure 1, and the full set of categories can be found in Appendix C.
@@ 3 Models
We begin by describing the CVCL (Child’s View for Contrastive Learning) model from Vong et al., (2024) before describing the variants we explore in this work. As described in the previous section, the input provided to the model is in the form of pairs of video frames and child-directed utterances. From this input, we need a method to align the representations from image frames with the representations of corresponding utterances, thereby linking visual referents to their corresponding words. To accomplish this feat, CVCL trains two neural networks jointly, learning to match image frames with their corresponding utterances. One of these networks (the image encoder) embeds video frames, while the other network (the language encoder), separately embeds utterances, projecting both into a shared multimodal representation space. This operationalizes the cross-situational learning approach from cognitive science Yu and Smith, (2007) through the lens of modern neural networks.
Specifically, CVCL’s vision encoder is designed to take in static video frames, passing them through a ResNeXt CNN architecture (Xie et al.,, 2016) that was pre-trained solely from the visual input from this child, and frozen except for a learnable linear projection head, to obtain a single embedding for each frame. CVCL’s language encoder is designed to take in transcribed child-directed utterances, passing each word through a single Embedding layer separately and then averaging across each word’s embedding in the utterance, again to obtain a single embedding for each utterance. Both encoders are jointly trained by passing embeddings into a contrastive loss, bringing closer matched embeddings from paired video frames and utterances, while separating mismatched embeddings from different video frames and utterances, similar to CLIP (Radford et al.,, 2021). This process of joint associative and representation learning produces a shared multimodal representation space, where neural similarity between words and referents determines the strength of their mapping.
We systematically explore three different architectural modifications to CVCL to examine their effects on grounded word learning. Each of these variants utilize the same underlying vision encoder architecture, but vary in the choice of language encoder and learning objective, each of which is illustrated in Figure  2.

### v2

@@ Abstract
What insights can machine learning bring to understanding human language acquisition? Large language and multimodal models have achieved remarkable capabilities, but their reliance on massive training datasets creates a fundamental mismatch with children, who succeed in acquiring language from comparatively limited input. To help bridge this gap, researchers have increasingly trained neural networks using data similar in quantity and quality to children’s input. Taking this approach to the limit, Vong et al. (2024) showed that a multimodal neural network trained on 61 hours of visual and linguistic input extracted from just one child’s developmental experience could acquire word-referent mappings. However, whether this approach’s success reflects the idiosyncrasies of a single child’s experience, or whether it would show consistent and robust learning patterns across multiple children’s experiences was not explored. In this article, we applied automated speech transcription methods to the entirety of the SAYCam dataset, consisting of over 500 hours of video data spread across all three children. Using these automated transcriptions, we generated multi-modal vision-and-language datasets for both training and evaluation, and explored a range of neural network configurations to examine the robustness of simulated word learning. Our findings demonstrate that networks trained on automatically transcribed data from each child can acquire word-referent mappings, generalizing across videos, children, and image domains. These results validate the robustness of multimodal neural networks for grounded word learning, while highlighting the individual differences that emerge in how models learn when trained on each child’s developmental experiences.
@@ 1 Introduction
Recent progress in large language models and multi-modal AI systems have led to fluent language generation and accurate object recognition, among other human-like abilities. These developments have naturally piqued the interest of many cognitive scientists and developmental psychologists, raising the question of whether these computational advances can help explain how children acquire language in the first few years of their lives. Yet, there is a striking asymmetry to reckon with: can machines trained to learn language from billions, even trillions, of word tokens provide insight into how children learn language from mere millions (Frank, 2023)? Addressing this disparity requires examining whether these language learning mechanisms can succeed under more realistic developmental constraints, combining these computational methods with richer and more ecologically valid datasets from developmental psychology.
@@
Recent efforts towards this goal have examined what can be learned from linguistic (and other sensory) data that is comparable in quantity or kind to what children receive naturally (Huebner et al., 2021; Wang et al., 2023; Warstadt et al., 2024; Qin et al., 2023; Vong et al., 2024; Hu et al., 2024; Zhuang et al., 2024a). For example, Huebner et al. (2021) examined language learning in neural networks from aggregated child-directed speech corpora (MacWhinney, 1992), demonstrating that models can acquire syntactic and semantic knowledge in the limited-data regime that is more representative of the kinds of input children receive. Similarly, the BabyLM challenge (Warstadt et al., 2024) examined what kinds of linguistic competencies can be learned from training language models on datasets consisting of 10M to 100M words (Hu et al., 2024; Charpentier et al., 2025). Separately, Zhuang et al. (2024a) studied how visual information can improve language learning, especially when using developmentally plausible amounts of data.
@@
Additional recent work has aimed to increase the realism of the training data by using egocentric, head-mounted camera recordings from developing children, such as SAYCam (Sullivan et al., 2022) and BabyView (Long et al., 2024). These datasets provide a unique glimpse into what children actually see and hear over the course of language learning, providing a testbed for what can be learned by training models from this kind of naturalistic, high-quality input. Researchers have shown that certain kinds of syntax and semantics, as well as word-referent mappings, are all learnable, even under the extreme constraint of only using data collected from a single child (Wang et al., 2023; Qin et al., 2023; Vong et al., 2024; Long et al., 2024).
@@
In recent work most related to the current article, Vong et al. (2024) demonstrated that a multimodal neural network trained on vision-and-language data extracted from a single child could acquire word-referent mappings. Specifically, they used a subset of a single child in SAYCam (Sullivan et al., 2022), curating a dataset of paired child-directed utterances with corresponding video frames from 61 hours of video. This dataset was used to train a multimodal neural network utilizing two separate encoders for visual and linguistic inputs, combined with a contrastive objective to align embeddings from both modalities into a shared multimodal representation space. The model’s knowledge of word-referent mappings was evaluated on two tasks involving 22 within-distribution categories and 64 out-of-distribution categories, achieving high performance on the former and modest performance on the latter. This work demonstrated that grounded word-referent mappings could be acquired from a subset of one child’s experience from relatively generic neural networks. However, a major limitation is that its findings are reliant on the data from a single child (as only one child’s data had transcripts at the time), raising questions about how robust or generalizable the findings are.
@@
In this article, we examine the robustness of the findings reported in Vong et al. (2024) by utilizing the full 500 hours of video from the SAYCam dataset (Sullivan et al., 2022) and training distinct models on each child’s data. The main contributions of our study are:
@@
  • We examine how learned word-referent mappings generalize beyond training data, finding successful generalization across videos, across children, and across image domains, with some degree of transfer even under substantial domain shifts.
  • We explore robustness across different language model configurations and training objectives, finding modest variations in performance, suggesting that dataset characteristics outweigh specific modeling or architectural choices.
@@ 2 Datasets
Our approach requires datasets that closely mirror the visual and linguistic experience from the perspective of developing infants. One dataset that fits this criterion is SAYCam (Sullivan et al., 2022), a video dataset capturing the egocentric, longitudinal perspective of 3 different children aged 6 to 32 months, collected via head-mounted cameras worn for approximately two hours per week. Parents were instructed to record naturalistically during everyday activities, resulting in footage covering mealtimes, play, reading, and outdoor activities. The three children (S - male, A - female, Y - male) were all from English-speaking families in the United States and Australia, each with a primary caregiver who was a psychologist. One child (S) was diagnosed with autism spectrum disorder at age 3, after the recording period; as of age 7, S was fully mainstreamed and did not require any special support. The other two children were typically developing.
While previous attempts to train language and multimodal models from SAYCam showed promise in answering questions around learnability (Wang et al., 2023; Vong et al., 2024; Qin et al., 2023), their results were limited by the availability of manually transcribed natural language data, primarily to a subset of one of the three children’s data. In this work, we expand the amount of available data for training by a factor of 7 through the use of automated speech transcription tools (Radford et al., 2023; Bain et al., 2023). This enabled the creation of new training and evaluation datasets for each of the three children in SAYCam, providing an opportunity to replicate our studies of simulated word learning with multiple children’s data.
@@ 2.1 Training Datasets
Transcribing videos. In order to train separate vision-language models on each child’s data, we needed a way to transcribe all of the speech from the raw videos. To efficiently transcribe the long-form audio from 500 hours of video from SAYCam, we used WhisperX (Bain et al., 2023), a tool which leverages the transcription capabilities of OpenAI’s Whisper model (Radford et al., 2023) (using the v3-large model with the language set to English), and applies further processing for efficient transcription and speaker diarization and word-level timestamps. This provided us with the transcribed text of each separate utterance, the start and end timestamps, and the predicted speaker label. Additional details surrounding the automatic transcription procedure and evaluation for speech transcription accuracy can be found in Zhang et al. (2025).
@@
Due to the inherent challenges of recording audio from a child’s head-mounted camera, the speech intelligibility can vary greatly across contexts, especially for child-directed speech. Nevertheless, an analysis comparing a subset of manually vs. automatically transcribed speech content showed that it was relatively accurate for adult speech and also sufficiently similar to the previously manually transcribed subset (Zhang et al., 2025). Furthermore, although the relative timings for the detected start and end timestamps also showed some discrepancies, this is not necessarily a major issue: Vong et al. (2024) observed that imperfect timing information (in the manual transcriptions) were still sufficient for models to acquire word-referent mappings.
@@
Whisper was not a panacea for annotating the data: it showed relatively poor performance on diarization accuracy (correctly distinguishing between different speakers in a video). Therefore, we did not attempt to distinguish between different speakers (e.g. child-directed versus child-produced speech), and retained all of the transcribed utterances. Another limitation was from hallucinations in the Whisper transcriptions, which resulted in instances of repetitive utterances, [^1] This included filtering repetitive phrases within a single utterance, as well as filtering adjacent utterances with the same content. In both cases, we only kept a single instance of the repeated utterance., requiring them to be filtered out. Additional pre-processing steps included removing all punctuation and lower-casing the transcribed text. Tokenization was performed at the word level using spaCy (Honnibal and Montani, 2017) and was subsequently used to create separate tokenizers from the transcribed vocabulary of each dataset split (after excluding any words with a frequency of two or fewer). After filtering and pre-processing, this resulted in 266K automatically transcribed utterances across the entire SAYCam dataset.
@@
Training datasets. To create vision-language datasets for training, we used this set of transcribed utterances to extract video frames in their original resolution that were temporally aligned with each utterance. For each detected utterance, we used the detected starting timestamp to extract up to 16 frames from the video at a rate of 3.75 fps, corresponding to 4.27s of video. Each set of extracted video frames was paired with their corresponding utterance, forming the basis of our vision-language dataset for training. The total amount of available training data from automating the transcription process is roughly 7 times larger than what was previously available for a single baby via manual transcription (Sullivan et al., 2022; Vong et al., 2024). We split the utterances into five different dataset splits, making sure that each split uses data from at most one child. The five different splits are:
@@
  • S-Whisper-2022: This split is based on the subset of the Whisper utterances from the same set of videos that were previously manually transcribed for baby S, as described in Sullivan et al. (2022), and used in Vong et al. (2024).
@@ 2.2 Evaluation Datasets
Figure 1: Evaluation frames from Labeled-S-V2, Labeled-A and Labeled-Y. Here we present three randomly selected evaluation frames from four different evaluation categories (ball, toy, book and cereal), with each row indicating frames derived from each of the three children in the SAYCam dataset.
For model evaluation, we primarily focus on image classification tasks, which are based on the looking-while-listening paradigm from early word learning studies (Golinkoff et al., 1992). In this evaluation procedure, models are presented with a target word along with four images (one belonging to the target category, and three distractors from other randomly sampled categories). The task is to select the image corresponding to the target word, where these labeled frames are sourced either from SAYCam directly, or other image datasets. Our model evaluations focus primarily on concrete nouns, as they are very frequent in training and straight-forward to construct evaluations for. [^2] Other lexical categories, while potentially learnable, pose additional evaluation challenges in egocentric videos. For example, verbs often co-occur with parents narrating their child’s actions, such that actions like “run” and “jump” are primarily associated with motion-blurred frames, while adjectives involving more abstract or relational properties are less reliably grounded directly in visual features. We test our models on a number of different evaluation datasets:
  • Labeled-S: This first evaluation set consists of 22 different visual categories derived from baby S’s egocentric video data, covering the most common kinds of objects and places both seen and spoken about (Orhan et al., 2020; Vong et al., 2024). For each category, there are 100 separate evaluation trials for a total of 2200 evaluation trials. The full set of categories can be found in Appendix D. This evaluation dataset is used to measure within-child generalization.
  • Labeled-S-V2, Labeled-A and Labeled-Y: These three additional evaluation datasets were generated using frames from each child’s own test split, enabling us to measure both within-child and cross-child generalization.  [^3] For Labeled-S-V2, due to some differences in dataset splitting, there was some partial overlap between the videos used for extracting frames for training and evaluation. These evaluation datasets were constructed using a mixture of automatic and manual filtering (see Appendix A for additional details), resulting in the same set of 41 object categories for each child’s evaluation dataset. [^4] The 41 categories partially overlap with, but are not a superset of, the original 22 Labeled-S categories. These 41 categories were selected based on common visual concepts present to all three children’s environments. Each evaluation dataset consists of 100 trials per category, for a total of 4100 evaluation trials. These 41 categories cover common visual concepts encountered by each child in SAYCam, including body parts (arm, hair, knee), clothing (shirt, jeans, shoe), living spaces (kitchen, porch, table), objects (ball, bottle, boy) and food (banana, cereal). A limited set of evaluation frames are shown in Figure 1, and the full set of categories can be found in Appendix D.
  • Konkle Objects: This evaluation dataset consists of 60 visual categories across 1,784 trials derived from naturalistic photographs of everyday objects on white backgrounds (Konkle et al., 2010), enabling us to measure out-of-distribution generalization. Similar to the above evaluation datasets, these 60 categories are also present in each of the children’s vocabularies, facilitating cross-dataset comparisons.
@@ 3 Models
We begin by describing the CVCL (Child’s View for Contrastive Learning) model from Vong et al. (2024) before describing the variants we explore in this work. As described in the previous section, the input provided to the model is in the form of pairs of video frames and child-directed utterances. From this input, we need a method to align the representations from image frames with the representations of corresponding utterances, thereby linking visual referents to their corresponding words. To accomplish this feat, CVCL trains two neural networks jointly, learning to match image frames with their corresponding utterances. One of these networks (the image encoder) embeds video frames, while the other network (the language encoder), separately embeds utterances, projecting both into a shared multimodal representation space. This operationalizes the cross-situational learning approach from cognitive science Yu and Smith (2007) through the lens of modern neural networks.
Specifically, CVCL’s vision encoder is designed to take in static video frames, passing them through a ResNeXt CNN architecture (Xie et al., 2016) that was pre-trained solely from the visual input from this child, and frozen except for a learnable linear projection head, to obtain a single embedding for each frame. CVCL’s language encoder is designed to take in transcribed child-directed utterances, passing each word through a single Embedding layer separately and then averaging across each word’s embedding in the utterance, again to obtain a single embedding for each utterance. Both encoders are jointly trained by passing embeddings into a contrastive loss, bringing closer matched embeddings from paired video frames and utterances, while separating mismatched embeddings from different video frames and utterances, similar to CLIP (Radford et al., 2021). This process of joint associative and representation learning produces a shared multimodal representation space, where neural similarity between words and referents determines the strength of their mapping.
We systematically explore three different architectural modifications to CVCL to examine their effects on grounded word learning. Each of these variants utilizes the same underlying vision encoder architecture, but varies in the choice of language encoder and learning objective, each of which is illustrated in Figure 2.